What is the net force on an airplane window of area 1000 cm2 if the pressure inside the cabin is 0.97 atm and the pressure outside is 0.85 atm? (Outward is defined as positive.)
Answer #1

IT 4- 4
Answer #2

SOLUTION :


Inside pressure = 0.97 atmosphere = 0.97 * 101325 = 98285.25  Pa


Outside pressure = 0.85 atmosphere = 0.85* 101325 = 86126.25 Pa 


These pressures act in opposite directions.


So, net pressure on the window = 98285.25 - 86126.25 = 12159 Pa = 12159 N/m^2


So, 


Force on the window 

= Net pressure  in N/m^2  * Area in m^2

= 12159 *  1000 / 100^2   ( N )

= 1215.9 N (ANSWER).

answered by: Tulsiram Garg